Can someone help me solve this please:
A boat starts in 15 minutes, and the landing is 1 mile away. If you can walk at 3 mph and you run at 6 mph, how many minutes may you walk and must you run to reach in time?

Answer by Alan3354(69443)

A boat starts in 15 minutes, and the landing is 1 mile away. If you can walk at 3 mph and you run at 6 mph, how many minutes may you walk and must you run to reach in time?
-------------------
To get there in 15 minutes, you have to average 4 mi/hr. (15 min = 1/4 hr)
----------
d = distance running at 6 mi/hr
d/6 + (1-d)/3 = 1/4
2d + 4-4d = 3
d = 1/2 mile @ 6 mi/hr
--> 1/2 mile at 6 mi/hr and 1/2 at 3 mi/hr to arv in 15 minutes